Jumper cables used to start a stalled vehicle often carry a 65-A current. How strong is the magnetic field 5.5 cm from one cable? Compare to the Earth's magnetic field (5.0×10−5 T).
Answer #1

here,

current , I = 65 A

distance from cable , r = 0.055 m

magnetic feild , B = u0*I /( 2 * pi * r)

B = 1.26 * 10^-6 * 65 /( 2 * pi * 0.055)

B = 2.36 * 10^-4 T

the magnetic feild is 2.36 * 10^-4 T

Be = 5 * 10^-5 T

B = 2.36 * 10^-4 /( 5 * 10^-5) * Be

B = 4.73 * Be

the magnetic feild is 4.73 times the earth's magnetic feild